# Mathematical Insights into Protein Architecture: Persistent Homology and Machine Learning Applied to the Flagellar Motor

### v2

@@ Abstract
organization=Ibn Tofail University, addressline=B.P. 242, city=Kenitra, postcode=14000, state=Kenitra, country=Morocco
@@
organization=ESISA, addressline=ESISA Analytica Laboratory, city=Fez, postcode=30050, state=Fez-Meknes, country=Morocco
@@
organization=Fordham University, addressline=Department of Computer and Information Science, Fordham University, Bronx, city=New York, postcode=10023, state=New York, country=USA
@@ 1 Introduction
Proteins are fundamental biological macromolecules that perform various functions in living organisms, including enzymatic catalysis, signal transduction, and structural support [1]. The three-dimensional shape of a protein determines its functionality, making the study of protein structure a cornerstone of molecular biology [2].
Recent advances in high-throughput technologies, such as X-ray crystallography, cryo-electron microscopy, and deep sequencing, have generated vast amounts of structural and functional data. However, analyzing and interpreting this data remains a significant challenge due to the complexity of protein structures and their dynamic nature [3][4][5].
The use of topological methods, particularly persistent homology, has gained traction in the study of protein structures [6]. Persistent homology has been applied to identify critical features of protein folding; Analyze binding sites and active regions of enzymes and understand dynamic conformational changes in proteins.
Moreover, machine learning approaches have been integrated with topological descriptors, demonstrating significant potential in predicting protein functionality and stability [7][8]. Despite these advances, the application of more refined tools, such as persistent homology computations, remains underexplored [9]. Although persistent homology effectively captures global and locomotive characteristics, it focuses primarily on presence of topological characteristics such as connected components, loops, and voids [6]. However, it often neglects additional algebraic structures, such as persistent homology classes, which provide richer invariants to understand interactions within protein structures [10].
Furthermore, there is limited exploration of how persistent homology computations can improve protein structure classification and provide new information on the relationships between structural and functional properties [11], as well as seamless integration with statistical and machine learning pipelines for predictive analysis.
@@ 2 Mathematical Model
By leveraging persistent homology, we gain access to a richer mathematical framework that transcends the limitations of traditional methods. It enables the identification of structural features tied to biological function, the study of conserved motifs, and the robust integration of data across varying scales. These strengths position persistent homology as a transformative tool for modern protein structure analysis. The significant contribution of this paper lies in providing an intrinsic framework for geometrical modeling of molecular shapes by investigating the high-dimensionality aspect provided by persistent homology, it is also recommendable to revisit our previous work [12][13][14][15], for more enlightenment about the topology function relationship paradigm of proteins, since this work is an extension to figure out the theoretical aspect of the homological degrees of topological representations, this will help also reduce dramatically the complexity of algorithms and provide an intrinsic framework for protein structure modeling.
@@
[…]
@@
[…]
@@
Each summand $\mathbb{k}[t]/(t^{n})$ corresponds to an interval $[0,n)$. Each free summand $\mathbb{k}[t]$ corresponds to $[0,\infty)$.
@@
This algebraic description captures both the topological changes and their lifetime across the filtration, using the language of modules over $\mathbb{k}[t]$
@@ 3 Implementation & Analysis
This section presents the implementation methodology using protein folding data from PDB (ID: 7CGO) [9].
@@
For a complete understanding of how to handle biomolecular data, the reader is referred to [16], [17], [18], [19], and [20].
@@
Kernel of $\text{Kernel of }\partial_{2}:\dim(\ker(\partial_{2}))=\text{len}(\ker_{2}),\quad\text{Image of }\partial_{1}:\dim(\text{im}(\partial_{1}))=\text{len}(im_{1})$ (8)
@@
The resulted matrix can subsequently be used as input for the persistent homology calculations. This provides a straightforward approach for extracting the shape of a protein, unlike traditional methods that use numerous complicated parameters to construct matrices intended to reconstruct the geometric conformation, as seen in molecular nonlinear dynamics and the flexibility-rigidity index involving exponential kernels with parameters. For a more detailed investigation into the relationship between topology and protein functions, we refer the reader to [21], [16], [22], and [23]. An interesting framework for understanding the computational aspects can also be found in [24], [25], [22], [26], [9], [27], and [28].
@@ 4.1 Data Collection & Feature Engineering
The dataset utilized in this study comprises 3D spatial configurations derived from Protein Data Bank (PDB) structures [29]. Specifically, the atomic coordinates (xyz) were extracted from .ent files downloaded from open-access repositories such as the RCSB Protein Data Bank [29]. These structures provide detailed molecular conformations of bacterial flagellar motors.
@@ 5 Comparison & Discussion
Smith et al. [3] Investigation of protein folding dynamics RMSD, temporal folding metrics Advanced molecular dynamics simulations Elucidated intermediate states and folding kinetics
Johnson et al. [7] Domain-specific protein classification Accuracy, precision, recall, F1 Convolutional neural architectures Achieved 92% accuracy via enhanced data augmentation
Lee et al. [8] Protein-protein interaction prediction AUC-ROC, precision-recall curves Graph neural networks with feature extraction State-of-the-art performance (AUC = 0.89)
Patel et al. [6] Topological protein characterization Topological invariants, persistence Persistent homology, barcode analysis Established topology-stability correlations
@@
Smith et al. [3] analyzed protein folding pathways using molecular dynamics simulations, tracking folding intermediates, and quantifying folding times. Their findings provided valuable insight into the kinetic and structural properties of protein folding.
Johnson et al. [7] explored protein domain classification through deep learning, employing convolutional neural networks (CNNs) alongside cross-validation and data augmentation. Their approach achieved an impressive accuracy of 92%, highlighting the robustness of deep learning techniques in domain prediction.
Lee et al. [8] investigated the prediction of protein-protein interaction using graph neural networks (GNNs). By extracting features from interaction data, their model achieved an area under the curve (AUC) score of 0.89, surpassing traditional prediction methods and setting a benchmark for future studies.
Patel et al. [6] applied persistent homology and barcode analysis to study the topological features of proteins. Their work established correlations between topological invariants, such as Betti numbers, and protein stability, revealing structural properties linked to folding behavior.

### v3

@@ Abstract
organization=Ibn Tofail University, addressline=B.P. 242, city=Kenitra, postcode=14000, state=Rabat-Salé-Kénitra, country=Morocco
@@
organization=Department of Mathematics, Faculty of Sciences and Techniques of Al Hoceima (FSTH), addressline=Abdelmalek Essaâdi University, city=Al Hoceima, postcode = 32003, state=Tanger-Tétouan-Al Hoceïma, country=Morocco
@@
organization=Department of Computer and Information Science, Fordham University, addressline=Bronx, city=New York, postcode=10023, state=New York, country=USA
[…]
organization=Centre Régional des Métiers de l’Éducation et de la Formation (CRMEF), addressline=Av. Allal Al Fassi, Rabat 11000, city=Rabat, postcode=10000, state=Rabat-Salé-Kénitra, country=Morocco
@@ 1 Introduction
Proteins are fundamental biological macromolecules that perform various functions in living organisms, including enzymatic catalysis, signal transduction, and structural support [b26]. The three-dimensional shape of a protein determines its functionality, making the study of protein structure a cornerstone of molecular biology [b27].
Recent advances in high-throughput technologies, such as X-ray crystallography, cryo-electron microscopy, and deep sequencing, have generated vast amounts of structural and functional data. However, analyzing and interpreting this data remains a significant challenge due to the complexity of protein structures and their dynamic nature [b23][b30][b31].
The use of topological methods, particularly persistent homology, has gained traction in the study of protein structures [b25]. Persistent homology has been applied to identify critical features of protein folding; Analyze binding sites and active regions of enzymes and understand dynamic conformational changes in proteins.
Moreover, machine learning approaches have been integrated with topological descriptors, demonstrating significant potential in predicting protein functionality and stability [b22][b24]. Despite these advances, the application of more refined tools, such as persistent homology computations, remains underexplored [b1]. Although persistent homology effectively captures global and locomotive characteristics, it focuses primarily on presence of topological characteristics such as connected components, loops, and voids [b25]. However, it often neglects additional algebraic structures, such as persistent homology classes, which provide richer invariants to understand interactions within protein structures [b29].
Furthermore, there is limited exploration of how persistent homology computations can improve protein structure classification and provide new information on the relationships between structural and functional properties [b28], as well as seamless integration with statistical and machine learning pipelines for predictive analysis.
@@ 2 Mathematical Model
By leveraging persistent homology, we gain access to a richer mathematical framework that transcends the limitations of traditional methods. It enables the identification of structural features tied to biological function, the study of conserved motifs, and the robust integration of data across varying scales. These strengths position persistent homology as a transformative tool for modern protein structure analysis. The significant contribution of this paper lies in providing an intrinsic framework for geometrical modeling of molecular shapes by investigating the high-dimensionality aspect provided by persistent homology, it is also recommendable to revisit our previous work [b18][b19][b20][b21], for more enlightenment about the topology function relationship paradigm of proteins, since this work is an extension to figure out the theoretical aspect of the homological degrees of topological representations, this will help also reduce dramatically the complexity of algorithms and provide an intrinsic framework for protein structure modeling.
@@
$C_{0}$ $=\langle A,B,C,D,E,F\rangle\quad\text{(degree 0)}$
$C_{1}$ $=\langle AB,AC,BC,DE,DF,EF\rangle\quad\text{(degree 1)}$
$C_{2}$ $=\langle ABC,DEF\rangle\quad\text{(degree 2)}$
@@
$\partial_{2}(\sigma)$ $=\text{sum of boundary edges of 2-simplex }\sigma$
@@
Each summand $\mathbb{k}[t]/(t^{n})$ corresponds to an interval $[0,n)$. Each free summand $\mathbb{k}[t]$ corresponds to $0,\infty)$.
@@
This algebraic description captures both the topological changes and their lifetime across the filtration, using the language of modules over $\mathbb{k}[t]$ .
@@ 3 Implementation & Analysis
This section presents the implementation methodology using protein folding data from PDB (ID: 7CGO) [b1].
@@
For a complete understanding of how to handle biomolecular data, the reader is referred to [b6], [b8], [b9], [b11], and [b12].
@@
$\text{Kernel of }\partial_{2}:\dim(\ker(\partial_{2}))=\text{len}(\ker_{2}),\quad\text{Image of }\partial_{1}:\dim(\text{im}(\partial_{1}))=\text{len}(im_{1})$ (8)
@@
The resulted matrix can subsequently be used as input for the persistent homology calculations. This provides a straightforward approach for extracting the shape of a protein, unlike traditional methods that use numerous complicated parameters to construct matrices intended to reconstruct the geometric conformation, as seen in molecular nonlinear dynamics and the flexibility-rigidity index involving exponential kernels with parameters. For a more detailed investigation into the relationship between topology and protein functions, we refer the reader to [b7], [b6], [b16], and [b15]. An interesting framework for understanding the computational aspects can also be found in [b2], [b3], [b16], [b0], [b1], [b4], and [b5].
@@ 4.1 Data Collection & Feature Engineering
The dataset utilized in this study comprises 3D spatial configurations derived from Protein Data Bank (PDB) structures [rcsb2024]. Specifically, the atomic coordinates (xyz) were extracted from .ent files downloaded from open-access repositories such as the RCSB Protein Data Bank [rcsb2024]. These structures provide detailed molecular conformations of bacterial flagellar motors.
@@ 5 Comparison & Discussion
Smith et al. [b23] Investigation of protein folding dynamics RMSD, temporal folding metrics Advanced molecular dynamics simulations Elucidated intermediate states and folding kinetics
Johnson et al. [b22] Domain-specific protein classification Accuracy, precision, recall, F1 Convolutional neural architectures Achieved 92% accuracy via enhanced data augmentation
Lee et al. [b24] Protein-protein interaction prediction AUC-ROC, precision-recall curves Graph neural networks with feature extraction State-of-the-art performance (AUC = 0.89)
Patel et al. [b25] Topological protein characterization Topological invariants, persistence Persistent homology, barcode analysis Established topology-stability correlations
@@
Smith et al. [b23] analyzed protein folding pathways using molecular dynamics simulations, tracking folding intermediates, and quantifying folding times. Their findings provided valuable insight into the kinetic and structural properties of protein folding.
Johnson et al. [b22] explored protein domain classification through deep learning, employing convolutional neural networks (CNNs) alongside cross-validation and data augmentation. Their approach achieved an impressive accuracy of 92%, highlighting the robustness of deep learning techniques in domain prediction.
Lee et al. [b24] investigated the prediction of protein-protein interaction using graph neural networks (GNNs). By extracting features from interaction data, their model achieved an area under the curve (AUC) score of 0.89, surpassing traditional prediction methods and setting a benchmark for future studies.
Patel et al. [b25] applied persistent homology and barcode analysis to study the topological features of proteins. Their work established correlations between topological invariants, such as Betti numbers, and protein stability, revealing structural properties linked to folding behavior.